1. Consider the K-Maps given below. For each K- Map 

i. Write the appropriate standard form (SOP/POS) of Boolean expression.

ii. Design the circuit using AND, NOT and OR gates.

iii. Design the circuit only by using 

• NAND gates if the standard form obtained in part (i) is SOP.

• NOR gates if the standard form obtained in pat (i) is POS.



1
Expert's answer
2021-06-04T13:28:42-0400


Part a) is for the left K-map, Part b) is for the right K-map.

a.

i)

"SOP=A'BC+ABC+AB'C'"

"POS=(A'B'+C')(A'B'+C)(A'B+C')(AB+C')(AB'+C)"


"(A'B'+C')=(A'+C')(B'+C')=(A'+C'+B')(A'+C'+B)\times"

"\times (B'+C'+A')(B'+C'+A)=(A'+C'+B')(A'+C'+B))(B'+C'+A)"


"(A'B'+C)=(A'+C)(B'+C)=(A'+C+B)(A'+C+B')\times"

"\times(B'+C+A)(B'+C+A')=(A'+C+B)(A'+C+B')(B'+C+A)"


"(A'B+C')=(A'+C')(B+C')=(A'+C'+B)(A'+C'+B')\times"

"\times(B+C'+A)(B+C'+A')=(A'+C'+B)(A'+C'+B')(B+C'+A)"


"(AB+C')=(A+C')(B+C')=(A+C'+B)(A+C'+B')\times"

"\times(B+C'+A)(B+C'+A')=(A+C'+B)(A+C'+B')(B+C'+A')"


"(AB'+C)=(A+C)(B'+C)=(A+C+B)(A+C+B')\times"

"\times(B'+C+A)(B'+C+A')=(A+C+B')(A+C+B')(B'+C+A')"


Standard form:

"POS=(A'+C'+B')(A'+C'+B))(B'+C'+A)(A'+C+B)(A'+C+B')\times"

"\times(B'+C+A)(B+C'+A)"


ii)



iii)

OR gate using NAND gate:




AND gate using NAND gate:



Then:




b)

I)

"SOP=A'B'C'D'+A'B'CD'+A'BC'D+A'BCD'+ABC'D'+"

"+ABC'D+ABCD+AB'C'D'+AB'C'D+AB'CD+AB'CD'"


"POS=(A'B'+C'D)(A'B'+CD)(A'B+C'D')(A'B+CD)(AB+CD')"


ii)